# On the Potential of Microtubules for Scalable Quantum Computation

### v1

@@ VI Conclusions and Outlook
By integrating QED-based models of microtubule dynamics [3, 1, 4] with core principles of quantum information theory and using best available values for the salient physicochemical parameters, we conclude that our models predict that microtubules can function as multi-level qudit processors, contingent on the controllable preparation of coherent quantum states in tubulin dimers (methods for which are discussed elsewhere [33]).
In the current article, building upon en earlier representation [7] of classical MT tubulin dipole moments as a pseudospin non-linear $\sigma$-model, which admits various solitonic solutions, ranging from kinks and spikes to snoidal and helicoidal snoidal (double helix-like) non-linear wave solutions of dipole excitations, we have argued that these solitons can be viewed as coherent quantum states (or better pointer states) of tubulin dimer-dipole quanta (in the case of pointer states, the latter have not yet undergone complete collapse).
In this context we have identified the fundamental hexagonal unit of the honeycomb lattice representing the various tubulin heterodimer dipole arrangements in an MT (see fig. 4), with the fundamental unit of information storage, the so-called quDit, given that in our construction it is not the $\alpha$, $\beta$ tubulin dipole conformations that constitute the fundamental superimposed quantum states, but rather combinations of four dipole quantum states corresponding to the four corners of the parallelogram comprised of four out of the seven tubulin dimers that constitute the hexagonal fundamental cell (see fig. 5, parallelogram 0123 including the central tubulin dimer (0) and its neighboring dimers labelled (1,2,3)).
Upon the action of an external stimulus, the system of dimers in this fundamental parallelogram quantum fluctuates and is entangled (‘quantum wired’). Within the decoherence time of the tubulin dimers in the MT, the fundamental system makes a ‘decision’ on the optima path for signal and information/energy transduction/transport in an optimal, most efficient way, by means of the formation of appropriate solitonic states, which carry energy and signal in a dissipation-free manner. In this way, for instance, the double-helix snoidal waves are formed after the (incomplete) collapse of the various tubulin dimer quanta to appropriate pointer states [16], and they transport energy and signal in a dissipation-free way along the MT and even across the MT networks.
It is amusing to notice that snoidal waves, which here are used as information transporting entities in the MT network, appear also in the mathematical description of the dynamical equations of the pendulum. The latter provides a mechanical tool to keep track of the time. There are other ways of manipulating the information appropriately to tell the time, for instance either through definite sounds or by reading numbers on the table or the tower (turret) clocks. In particular, the turret clock for centuries was based on the pendulum, with its hands moving by means of an appropriate connection to hanging weights (pendulum). A similar role, as transporters of the passing of time, characterizes the snoidal solitonic waves of the dipole quanta in MT. There is a really good analogy between MT and turret clocks. The underlying mathematics (snoidal waves) are common to both systems. The snoidal waves are solutions of the pendulum equations of motion and of the equations of the MT systems, while the aforementioned conjunction of the turret clock hands to pendulum-type hanging weights also finds a nice analogy with the MAPs connecting the various MTs in a network (cf. fig. 7), which as we have discussed above and in [1] can also imply the operation of MT networks as logic gates (classical or quantum).
@@
In the present article we have discussed a QED cavity model for MT [3, 4], in which the ordered water interior of the MT operates as a highly isolated cavity, the interior walls of which are comprised by the tubulin dipole quanta. The strong interactions between the dimer dipole quanta and the dipole quanta of the ordered water near the walls imply a high-quality isolated cavity, with the main source of environmental entanglement being the leakage of water dipole quanta through the imperfect tubulin-protein-dimer walls of the MT cavity. As discussed in [3] and in the text above, this provides relatively long decoherence times of order $10^{-6}$ s, during which the entire MT network can be quantum wired (entangled) and energy and signal can be transported along a moderately long (of one micron long) MT.
We remark at this point that, in the absence of the cavity mechanism due to the strong ordered-water-dimer dipole-dipole interactions, the decoherence time of each tubulin dimer, as a consequence of the string dipole interaction with its neighboring dimers, would be significantly smaller, of order in the range ${\mathcal{O}}(1-100)$ fs (see (29)). Although the latter time suffices for quantum wiring of distances up to $40$ Angströms, which is appropriate for entangling the entire algae antennae [21, 20], unfortunately is not sufficient for quantum wiring of the entire MT or MT networks. Distances of that size span only the basic heagonal cell of tubulin dimer lattice (see fig. 5). To entangle the entire MT (or even beyond, the MT networks) one needs the QED cavity model which enhances the decoherence times of the dipole quanta of the heterodimers, as discussed in the previous text,
With the above mechanism we have completed our description of the basic requirements for a potential biocomputation mechanism involving MT networks, at ambient temperatures, as outlined in the beginning of section IV [50]. Specifically, as has been requested by repeated criticism of other publications in the field, we here designate explicitly: (i) a precise description of the quantum states of the quDit, the basic storage of information unit, in our case having a dimension $D=4$, in the simplest scenario (ii) a description of the mechanism through which the wavefunctions representing these states become entangled, including specification of the basis in which measurements of the quDit properties are performed in situ, and (iii) a detailed way of attaining quantum coherence over the required time scale so as to quantum-wire the MT efficiently. QuDit-gate operations in this framework are realized via dipole–dipole and dipole–field interactions, while solitonic excitations offer a plausible mechanism for coherent, low-loss quantum state transfer. This framework provides a biologically grounded platform for investigating quantum information processing in biotechnological, experimental settings, as the one proposed in Figure 7, and we propose that doing so can be used in advancing the prospect of scalable quantum computing.
For convenience we have assembled tables 1 and 2, where typical values of parameters that are relevant to proposed quantum (computation) values in biosystems, to be compared with the pertinent parameters of our cavity model of MT reviewed here. Using MTs as the substrate for biocomputers can only be resolved experimentally by performing experiments such as those proposed here and elsewhere in the literature. Another potential path is to attempt manmade quantum computers following the example of the MT, but using appropriate spin systems to construct the basic structures.
Affaire à suivre….
@@ Acknowledgments
The work of N.E.M. is supported in part by the UK Science and Technology Facilities research Council (STFC) under the research grant ST/X000753/1. NEM also acknowledges participation in the COST Association Actions CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)” and CA23130 “Bridging high and low energies in search of quantum gravity (BridgeQG)”. AM wishes to acknowledge the support received from www.RealNose.ai and the intellectual boost received from participants of the www.Osmocosm.Org MIT conferences and the MIT IAP class “Making Sense of Scent”.

### v2

@@ VI Conclusions and Outlook
By integrating QED-based models of microtubule (MT) dynamics [3, 1, 4] with principles of quantum information theory, and by employing the best available physicochemical parameters, we conclude that MTs can, in principle, function as multi-level qudit processors, provided that coherent quantum states of tubulin dimers can be controllably prepared (see [33]).
Building upon earlier representations [7] of classical MT tubulin dipole moments as a pseudospin non-linear $\sigma$-model, which admits solitonic solutions ranging from kinks and spikes to snoidal and helicoidal waves, we have argued that these solitons can be interpreted as coherent quantum pointer states of tubulin dimer-dipole excitations. In this construction, the fundamental hexagonal unit of the honeycomb lattice representing tubulin dipole arrangements (see Fig. 4) corresponds to the unit of quantum information storage, explicitly: a quDit. Unlike conventional approaches, in our model the quDit is not defined by the $\alpha$ and $\beta$ dipole conformations alone, but by combinations of four dipole quantum states associated with the parallelogram formed by four of the seven dimers in the fundamental hexagonal cell (see Fig. 5).
External stimuli drive quantum fluctuations and entanglement in this fundamental unit, leading—within the decoherence time of the MT—to a “decision” on optimal pathways for signal and energy transport via the formation of solitonic states. Double-helix snoidal waves, for instance, arise from the incomplete collapse of tubulin dimer quanta into pointer states [16] and propagate information along individual MTs and across MT networks in a dissipation-free manner. [^10] It is worth noting that snoidal waves, here proposed as information carriers in MTs, also emerge in the dynamics of the pendulum. This parallel has long underpinned timekeeping devices, such as turret clocks, which translate pendular oscillations into the motion of clock hands. In analogy, snoidal solitons in MTs convey the passage of time and information within cellular networks. The shared mathematics of pendulum motion and MT dynamics underscores this analogy, while the connections between MTs mediated by MAPs (cf. Fig. 8) echo the mechanical linkages that allow turret clocks to operate as logic devices.
@@
We have also discussed a QED cavity model for MTs [3, 4], wherein the ordered water in the lumen acts as a high-quality electromagnetic cavity, bounded by tubulin dipole quanta. Strong interactions between the dimer dipoles and the ordered-water dipoles near the walls yield a highly isolated cavity, with environmental coupling primarily due to leakage of water dipole quanta through the imperfect protein walls. As shown in [3], this configuration supports decoherence times on the order of $10^{-6}$ s, allowing micron-scale MTs to remain quantum-wired (entangled) and capable of dissipation-free energy and signal transport.
Absent this cavity effect, decoherence times for individual tubulin dimers—limited by direct dipole–dipole interactions—would be much shorter, $\mathcal{O}(1$$100)$ fs (see Eq. (29)). While such times suffice for quantum wiring across $\sim 40$ Å (adequate for entire photosynthetic antenna complexes [21, 20]), they are insufficient for coherent wiring of full MTs or MT networks. Thus, the cavity mechanism is essential to extend coherence times to biologically relevant scales.
Taken together, these considerations outline the requirements for biocomputation via MT networks at ambient temperatures, as introduced in Sec. IV [50]. Specifically, we have provided: (i) a precise description of the quDit states (dimension $D=4$ in the simplest case); (ii) a mechanism for entanglement of these states, including the measurement basis; and (iii) a pathway for maintaining coherence on the required timescale to support quantum wiring.
Within this framework, quDit-gate operations arise naturally from dipole–dipole and dipole–field interactions, while solitonic excitations enable coherent, low-loss state transfer. MT networks, therefore, present a biologically plausible platform for quantum information processing (Fig. 8), and their experimental interrogation could advance prospects for scalable quantum computation.
For reference, Tables 3 and 4 summarize typical parameter values relevant to proposed quantum information mechanisms in biosystems, compared with those in our MT cavity model. Ultimately, determining whether MTs can serve as substrates for biocomputation requires targeted experiments of the type proposed here and in related studies. A complementary approach is to design synthetic quantum devices inspired by MT architectures, for example by using engineered spin systems to replicate their essential features.
Before closing, we would like to mention several experimental directions. There are research works dealing with electron transport modeling in MT [94]. There are also studies on the role of MT as regulators of the shape and functions of network of neurons in the brain [95], as well as research works on the electrical behaviour of (bovine) MT (viewed as complex electrical networks) and its potential connection with the role of MT as a medium for evolutionary computation [96]. Moreover, there have been analyses on classical computing modellng of MT networks, attempting to understand potential links with their physiology, in particular reproduction of features related to degenerating neurons in disease states, such as Alzheimer’s disease  [97]. There have also been efforts aimed at the construction of an electrical analogue computer out of MT protofilaments [98], where using operational amplifiers, capacitors, and resistors, the authors designed analytically the bioelectronic circuit of the MT protofilament, in an attempt to understand computational aspects of these biological entities. Information transfer and storage in brain has also been studied from an MT perspective [99], considering them as a communication channel. Although the above works deal with classical computational aspects of MTs, nonetheless we may think of combining such attempts with our studies here, which make use of the quantum computing aspects of MT, viewing them as QED cavities.
Affaire à suivre…
@@ Acknowledgments
The work of N.E.M. is supported in part by the UK Science and Technology Facilities research Council (STFC) under the research grant ST/X000753/1. NEM also acknowledges participation in the COST Association Actions CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)” and CA23130 “Bridging high and low energies in search of quantum gravity (BridgeQG)”. AM wishes to acknowledge the support received from www.RealNose.ai and the intellectual boost received from participants of the www.OsmoCosm.org MIT conferences and the MIT IAP class “Making Sense of Scent”. DVN would like to thank his family Olga and Odysseas for encouragement and patience and the Digital Health Literacy & Policy Hub Foundation (digitalhealth-hub.com) for support.